OpenStudy (lgbasallote):

\[x(x^2 +y^2)^2(ydx - xdy) + y^6dy = 0\] anyone have any ideas?

OpenStudy (compassionate):

You want me too solve it?

OpenStudy (lgbasallote):

sure...if you can lol

OpenStudy (compassionate):

\[x(x^2 + y^2)^2 (0) = -dy^7\]

OpenStudy (anonymous):

divide it by y^6...it gets in to standard form

OpenStudy (lgbasallote):

0??/ o.O

OpenStudy (anonymous):

hint : (ydx - xdy)/y^2 = d(x/y)

OpenStudy (lgbasallote):

idk if that made sense to me o.O

OpenStudy (anonymous):

cmon divide the entire thing by y^6

OpenStudy (anonymous):

(x/y)*(1+(x/y)^2))^2 * d(x/y) + dy = 0

OpenStudy (anonymous):

put x/y = t

OpenStudy (lgbasallote):

wait...how did ydx - xdy)/y^2 = d(x/y) first?

OpenStudy (anonymous):

ok wat's d(x/y)

OpenStudy (lgbasallote):

dx/y?

OpenStudy (lgbasallote):

or do you mean derivative of 9x/y)?

OpenStudy (lgbasallote):

derivative of (x/y)*

OpenStudy (anonymous):

d(x/y) yep derivative of x/y

OpenStudy (anonymous):

What's your question?

OpenStudy (lgbasallote):

the one in the blue box @beeqay

OpenStudy (lgbasallote):

@A.Avinash_Goutham with respect to what??

OpenStudy (anonymous):

I mean you have an equation fine, but what is the question, find dy/dx find dx/dy... solve for y.. Solve for x? Possibilities are a countless, so what's your question?

OpenStudy (lgbasallote):

uhh i thought all differential equations instruction is to find the solutioN?

OpenStudy (lgbasallote):

well anyway that's what's needed...a solution

OpenStudy (anonymous):

Ok, but still, you're question isn't specific enough, do you want dy/dx or dx/dy? there's a difference

OpenStudy (lgbasallote):

trivial question....

OpenStudy (anonymous):

well it's a standard integrallll :P um the solutiion is y a function of x :P

OpenStudy (lgbasallote):

^whatever that meant haha

OpenStudy (anonymous):

????

OpenStudy (lgbasallote):

so back to the question pls?

OpenStudy (anonymous):

well the qn is solve :P

OpenStudy (lgbasallote):

how?/

OpenStudy (anonymous):

t(1+t2)dt + dy = 0

OpenStudy (anonymous):

integrate that

OpenStudy (lgbasallote):

where did that come from?!!

OpenStudy (anonymous):

(1+t2)2 * sorry

OpenStudy (anonymous):

t = x/y

OpenStudy (lgbasallote):

where did t come from =_=

OpenStudy (lgbasallote):

t = x/y so ?

OpenStudy (compassionate):

This thread is hilarious.

OpenStudy (lgbasallote):

indeed

OpenStudy (compassionate):

Do you still not understand?

OpenStudy (lgbasallote):

yeah

OpenStudy (compassionate):

Factor first.

OpenStudy (lgbasallote):

how is dt = ydx - xdy

OpenStudy (lgbasallote):

factor what?

OpenStudy (anonymous):

dt is that upon y2

OpenStudy (compassionate):

Lol.

OpenStudy (compassionate):

This is thread of the year.

OpenStudy (lgbasallote):

lol why?

OpenStudy (compassionate):

Factor your squared equations to simplify some more before going through all the work.

OpenStudy (compassionate):

Lol. You're totally not getting it.

OpenStudy (anonymous):

factor?!

OpenStudy (lgbasallote):

im dyslexic :p lol

OpenStudy (compassionate):

Yes, Avin. You can factor in this particular one.

OpenStudy (compassionate):

I do it all the time. It makes it easier.

OpenStudy (lgbasallote):

i cannot see anything factorable o.O

OpenStudy (compassionate):

What is it you're trying to solve?

OpenStudy (anonymous):

um it's avi r alvin :|

OpenStudy (lgbasallote):

okay...i do NOT get either of your solutions =_=

OpenStudy (anonymous):

ok :P